# Discrete weakly stable distributions

### v1

@@ Abstract
A random variable $X$ is strictly stable if a sum of independent copies of $X$ has the same distribution as $X$ up to scaling, and is stable (or ‘weakly stable’) if the sum has the same distribution as $X$ up to both scaling and shifting. Steutel and Van Harn studied ‘discrete strict stability’ for random variables taking values in the non-negative integers, replacing scaling by the thinning operation. We extend this to ‘discrete (weak) stability’, where shifting is replaced by the addition of an independent Poisson random variable. The discrete stable distributions are ‘Poisson–delayed Sibuya’ distributions, which include the Poisson and Hermite distributions. Very similar results have been proved independently by Townes [17].
@@ 1.1 Definitions
A random variable $X$ is strictly stable if a sum of independent copies of $X$ has the same distribution as $X$ up to scaling, and is (weakly) stable if the sum has the same distribution as $X$ up to both scaling and shifting.
@@
We say that $X$ is stable (or weakly stable) if for all $n$ there exist constants $a_{n}\geq 0$ and $b_{n}$ such that
@@
The goal of this paper is to define and study a discrete equivalent of weak stability. To do this, we need a discrete equivalent of shifting a real-valued random variable $X$ by a constant $b\in\mathbb{R}$ to get $X+b$. We propose that a suitable equivalent is ‘Poisson shifting’. For $b\geq 0$, we can ‘Poisson shift’ a discrete random variable $X$ by adding an independent Poisson random variable with mean $b$. We will denote this $X\oplus b=X+Z$, where $Z\sim\operatorname{Po}(b)$ is independent of $X$. We will give more details about the Poisson shift – including arguments in favour of the Poisson shift as the natural ‘discrete shift’ – later in Subsection 2.1.
We can now state our proposed new definition of discrete (weak) stability.
@@
We say that $X$ is discrete stable (or discrete weakly stable) if for all $n$ there exist constants $a_{n}\in[0,1]$ and $b_{n}>0$ such that
@@ 1.2 Theorems
We recall the following two foundational results on usual non-discrete stability. With the exception of some special cases, probability density functions for stable distributions cannot be written down in closed form, but the characteristic functions $\phi_{X}(u)=\mathbb{E}\exp(\mathrm{i}uX)$ can be. For full details (especially the characterisations in terms of characteristic functions, which we omit), see, for example, [4, Chapters VI and XVII], [12, Chapters 1 and 3].
@@
If $X$ is strictly stable, in that (2) holds, and $X$ is not a point mass, then $a_{n}=n^{-1/\alpha}$ for some $\alpha\in(0,2]$.
@@
The main result of this paper, is the following result for discrete stability – discrete equivalent of Theorem 1.5.
@@
Independently of this work, the preprint “Broadly discrete stable distributions” by F. William Townes [17] covers the same topics as this paper and proves very similar results. When assigning credit for results, readers should note that it is Townes’s paper [17] that was publicly available first.
@@ 2.1 Properties of thinning and the Poisson shift
    $\operatorname{Disp}(X\oplus b)=\operatorname{Disp}(X)$ is the dispersion $\operatorname{Disp}(X)$ exists and is finite.
@@
    $\operatorname{Var}(X+b)=\operatorname{Var}(X)$ is the variance $\operatorname{Var}(X)$ exists and is finite.
@@ 2.2 Properties of the alternate probability generating function
We write $C(u)=\log\psi_{X}(u)$ for the factorial cumulant generating function. (We have taken the name from [8], although they actually apply that name to $\log\psi_{X}(-u)$.) If the APGF is the exponential of a complicated expression, this can be more convenient to deal with.
@@ 4 Proof of main theorem
We have proved that the Poisson–delayed Sibuya distributions have APGFs as in (5) and are discrete stable with $a_{n}=n^{-1/\alpha}$ for $\alpha\in(0,2]$. The remaining technical part of our main result, Theorem 1.7, is to show that $a_{n}$ must be of the form $a_{n}=n^{-1/alpha}$ for $\alpha\in(0,2]$, and that the APGF must be that of the Poisson–delayed Sibuya distribution as in (5).
@@ 4.1 Two useful lemmas
by continuity of $C$. So either $C(u)=\frac{b}{1-a}u$ for all $u\in[0,2]$, in which case $X$ is Poisson with rate $\frac{b}{1-a}$, which was forbidden in the statement of the lemma, or else $a=1$.
@@ 4.2 Showing that $a_{n}=n^{-1/\alpha}$
The proof that $\alpha$ must be in $(0,2]$ will come later, in Subsection [REF].
@@
Lemma 4.3 shows that $(a_{n})$ is completely multiplicative. If we can show that $(a_{n})$ is also decreasing, then that will prove that $a_{n}=n^{\beta}$ for some $\beta<1$ (see, for example, [5]), and we can set $\beta=-1/\alpha$.
@@ 4.3 Formulas for $b_{n}$
[…]
@@ 4.4 Form of the alternate probability generating function
$\mathbb{P}(Y=1)=-\psi^{\prime}Y(1)=-\frac{-\delta-\gamma}{\delta}=\frac{\delta+\gamma}{\delta}.$
Then $\psi^{\prime}Y(1)\geq 0$ requires $\delta+\gamma\geq 0$, or $\delta\geq-\gamma$, and $\psi^{\prime}_{Y}(1)\leq 1$ requires $\delta+\gamma\leq\delta$, or $\gamma\leq 0$. ∎
@@ Acknowledgements
The author thanks Sam Power for useful comments and information.

### v2

# Discrete broadly stable distributions

@@ Abstract
A random variable $X$ is strictly stable if a sum of independent copies of $X$ has the same distribution as $X$ up to scaling, and is stable (in the broad sense) if the sum has the same distribution as $X$ up to both scaling and shifting. Steutel and Van Harn studied ‘discrete strict stability’ for random variables taking values in the non-negative integers, replacing scaling by the thinning operation. We extend this to ‘discrete stability’ (in the broad sense), where shifting is replaced by the addition of an independent Poisson random variable. The discrete stable distributions are ‘Poisson–delayed Sibuya’ distributions, which include the Poisson and Hermite distributions. Similar results have been proved independently by Townes [17].
@@ 1.1 Definitions
A random variable $X$ is strictly stable if a sum of independent copies of $X$ has the same distribution as $X$ up to scaling, and is stable (in the broad sense) if the sum has the same distribution as $X$ up to both scaling and shifting.
@@
We say that $X$ is stable (in the broad sense) if for all $n$ there exist constants $a_{n}\geq 0$ and $b_{n}$ such that
@@
The goal of this paper is to define and study a discrete equivalent of stability in the broad sense. To do this, we need a discrete equivalent of shifting a real-valued random variable $X$ by a constant $b\in\mathbb{R}$ to get $X+b$. We propose that a suitable equivalent is ‘Poisson shifting’. For $b\geq 0$, we can ‘Poisson shift’ a discrete random variable $X$ by adding an independent Poisson random variable with mean $b$. We will denote this $X\oplus b=X+Z$, where $Z\sim\operatorname{Po}(b)$ is independent of $X$. We will give more details about the Poisson shift – including arguments in favour of the Poisson shift as the natural ‘discrete shift’ – later in Subsection 2.1.
We can now state our proposed new definition of discrete stability (in the broad sense).
@@
We say that $X$ is discrete stable (in the broad sense) if for all $n$ there exist constants $a_{n}\in[0,1]$ and $b_{n}>0$ such that
@@ 1.2 Theorems
We recall the following two foundational results on usual non-discrete stability. With the exception of some special cases, probability density functions for stable distributions cannot be written down in closed form, but the characteristic functions $\phi_{X}(t)=\mathbb{E}\exp(\mathrm{i}tX)$ can be. For full details (especially the characterisations in terms of characteristic functions, which we omit), see, for example, [4, Chapters VI and XVII], [12, Chapters 1 and 3].
@@
If $X$ is stable, in that (2) holds, and $X$ is not a point mass, then $a_{n}=n^{-1/\alpha}$ for some $\alpha\in(0,2]$.
The stable random variables can be fully characterised in terms of their characteristic functions $\phi_{X}$.
@@
The main result of this paper, is the following result for discrete stability: a discrete equivalent of Theorem 1.5.
@@
Independently of this work, the preprint “Broadly discrete stable distributions” by F. William Townes [17] covers the same topics as this paper and proves very similar results.
@@ 2.1 Properties of thinning and the Poisson shift
    $\operatorname{Disp}(X\oplus b)=\operatorname{Disp}(X)$ if the dispersion $\operatorname{Disp}(X)$ exists and is finite.
@@
    $\operatorname{Var}(X+b)=\operatorname{Var}(X)$ if the variance $\operatorname{Var}(X)$ exists and is finite.
@@ 2.2 Properties of the alternate probability generating function
We write $C(t)=\log\psi_{X}(t)$ for the factorial cumulant generating function. (We have taken the name from [8], although they actually apply that name to $\log\psi_{X}(-u)$.) If the APGF is the exponential of a complicated expression, this can be more convenient to deal with.
@@ 4 Proof of main theorem
We have proved that the Poisson–delayed Sibuya distributions have APGFs as in (5) and are discrete stable with $a_{n}=n^{-1/\alpha}$ for $\alpha\in(0,2]$. The remaining technical part of our main result, Theorem 1.7, is to show that $a_{n}$ must be of the form $a_{n}=n^{-1/\alpha}$ for $\alpha\in(0,2]$, and that the APGF must be that of the Poisson–delayed Sibuya distribution as in (5).
@@ 4.1 Two useful lemmas
by continuity of $C$. So either $C(u)=-\frac{b}{1-a}u$ for all $u\in[0,2]$, in which case $X$ is Poisson with rate $\frac{b}{1-a}$, which was forbidden in the statement of the lemma, or else $a=1$.
@@ 4.2 Showing that $a_{n}=n^{-1/\alpha}$
The proof that $\alpha$ must be in $(0,2]$ will come later, in Subsection 4.4.
@@
Lemma 4.3 shows that $(a_{n})$ is completely multiplicative. If we can show that $(a_{n})$ is also decreasing, then that will prove that $a_{n}=n^{\beta}$ (see, for example, [5]) for some $\beta<0$ , and we can set $\beta=-1/\alpha$.
@@ 4.3 Formulas for $b_{n}$
Part 1 is the same as the start of the proof for the $\alpha\neq 1$ case. We go back to (9), set $a_{n}=n^{-1}$, and get
@@ 4.4 Form of the alternate probability generating function
$\mathbb{P}(Y=1)=-\psi^{\prime}_{Y}(1)=-\frac{-\delta-\gamma}{\delta}=\frac{\delta+\gamma}{\delta}.$
Then $\psi^{\prime}_{Y}(1)\geq 0$ requires $\delta+\gamma\geq 0$, or $\delta\geq-\gamma$, and $\psi^{\prime}_{Y}(1)\leq 1$ requires $\delta+\gamma\leq\delta$, or $\gamma\leq 0$. ∎
@@ Acknowledgements
The author thanks Nadjib Bouzar, Sam Power, and Will Townes for useful discussions and information.